OpenStudy (anonymous):

A typical car depreciates in value by about 15% per year. How long would it take a $35000 car to be valued at $15000

OpenStudy (anonymous):

solve \[15,000=35000(.85)^t\] for t

OpenStudy (anonymous):

so subtract the 15000 from 35000? or?

OpenStudy (anonymous):

\[\frac{3}{7}=.85^t\] \[t=\frac{\ln(\frac{3}{7})}{\ln(.85)}\]

OpenStudy (anonymous):

first divide, then use change of base

OpenStudy (anonymous):

divide by 3500?

OpenStudy (anonymous):

yes divide by 35000

OpenStudy (anonymous):

wait, whats the formula you used?

OpenStudy (anonymous):

that is where i got the \[\frac{3}{7}\] from

OpenStudy (anonymous):

lets start at the beginning. the car loses 15% of its value every year, so every year it is worth 85% of the previous years value. that is how i got \[35,000\times (.85)^t\]

OpenStudy (anonymous):

now you want to know when it is worth 15,000 so put \[15,000=35,000\times (.85)^t\] and solve for t first step is to divide by 35,000 to get \[\frac{3}{7}=.85^t\] so far so good?

OpenStudy (anonymous):

yes. the whole lm thing confuses me

OpenStudy (anonymous):

just like in the next question you posted, so solve \[b^t=A\] for t use \[t=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

so if i see \[.85^t=\frac{3}{7}\] i solve for t in on step and write \[t=\frac{\ln(\frac{3}{7})}{\ln(.85)}\]

OpenStudy (anonymous):

thank you sooo much. so now how do i solve the other problem that i posted.